OpenStudy (anonymous):

don't understand can u help me What is the value of the expression All of 3.6 multiplied by 10 to the power 8 over all of 1.2 multiplied by 10 to the power 3? 2.4 × 105 3.0 × 1011 3.0 × 105 2.4 × 1011 -------------------------------------------------------------------------------- Question 4 (Multiple Choice Worth 5 points) (01.01 LC) Which rational number equals 0 point 3 with bar over 3? 1 over 5 1 over 3 3 over 10 3 over 5 --------------------------------------------------------------------------------

OpenStudy (anonymous):

Question 8 (Multiple Choice Worth 5 points) (01.05 LC) If x2 = 30, what is the value of x? ±60 ±15 ±square root of 30 ±square root of 15

OpenStudy (anonymous):

Did you finish any of them?

OpenStudy (anonymous):

I finished all the other questions on my test. but these im confused with

OpenStudy (anonymous):

and I need help

OpenStudy (anonymous):

The first one is actually \[\frac{ 3.6\times 10^{8} }{ 1.2\times 10^{3} }\] Can you tell the answer?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

For the first two numbers, all you need to do is dividing 3.6 by 1.2. For the exponent, the bases is identical (10), so you need to subtract the exponent of the top to the exponent of the bottom.

OpenStudy (anonymous):

oh ok thx <3

OpenStudy (anonymous):

can u help me with the rest

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

The second one is actually like this\[\frac{ ? }{ 3 }=0.3\] Can you find "?" ?

OpenStudy (anonymous):

I think it is 1

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Is the 3 numerator of denominator?

OpenStudy (anonymous):

denominator

OpenStudy (anonymous):

ok this one I was confused about the 0. There are (24)2 ⋅ 20 books in a cupboard. What is the total number of books in the cupboard?

OpenStudy (anonymous):

Sorry I cannot see the problem very clear. It gets lots of "?". Can you open a new thread and tag me in?